OpenStudy (anonymous):

solve the equation 3l2x+5l= 9x-6

OpenStudy (anonymous):

its solving absolute value equations, does anyone get this one?

OpenStudy (anonymous):

@hfenton Hahaha, sorry, sweety. I'm trying to remember, the last time I did absolute values was a long while ago. But no worries, I'm sure there are a lot of other members who know. Just be patient. And I hope your question gets answered. :)

OpenStudy (anonymous):

Start with dividing both sides for 3:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Please show your work right HERE!

OpenStudy (anonymous):

l2x+5l=3x-2

OpenStudy (anonymous):

Great :) Now in order to remove the absolute sign, we divide into 2 cases!

OpenStudy (anonymous):

First one is positive value: 2x +5 = 3x -2 Can you solve it?

OpenStudy (anonymous):

Second case is negative value: 2x + 5 = - ( 3x -2) Please show your work :)

OpenStudy (anonymous):

@hfenton Where have you gone???